OpenStudy (anonymous):

derivative of xe^(-2x)

OpenStudy (anonymous):

use product rule

OpenStudy (solomonzelman):

and chain rule for the exponent of -2x.

OpenStudy (anonymous):

is it e^(-2x)*(1-2x)

OpenStudy (fibonaccichick666):

i recommend using a substitution to make the comprehension a little easier

OpenStudy (fibonaccichick666):

how did you get that vintage type? Can you show us your work so we can guarentee every step is correct?

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{d}{dx}xe^{-2x} = 1\times e^{-2x}+xe^{-2x}\times(-2x) }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{d}{dx}xe^{-2x} = e^{-2x}-2x^2e^{-2x} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{d}{dx}xe^{-2x} = e^{-2x}(1-2x^2) }\) you were a bit off.

OpenStudy (some.random.cool.kid):

hi solomon

OpenStudy (anonymous):

\[\frac{ d }{ dx }\left( x e ^{-x} \right)=x*e ^{-2x}*(-2)+1*e ^{-2x}=\left( 1-2x \right)e ^{-2x}\]

OpenStudy (anonymous):

so you are correct vintage Typewriter.

OpenStudy (solomonzelman):

ophh true, I miss did it

OpenStudy (solomonzelman):

I didn;t derive the exponent, but just multiplied by it

OpenStudy (anonymous):

correction\[write \frac{ d }{ dx }\left( x e ^{-2x} =\right)\]